# NeuralFuse: Learning to Recover the Accuracy of Access-Limited Neural Network Inference in Low-Voltage Regimes

### v2

@@ Appendix I Additional Experiments on Adversarial Training
K-bits C.A. P.A. C.A. (NF) P.A. (NF) R.P. 100 92.6 38.9 12.4 92.4 38.3 12.1 -0.6 500 92.1 38.7 12.5 -0.2 5,000 92.6 38.9 12.5 0 20,000 60.1 23.0 8.1 -16 100,000 71.1 23.6 6.6 -16 • [Note] C.A. (%): clean accuracy, P.A. (%): perturbed accuracy, NF: NeuralFuse, and R.P.: total recover percentage of P.A. (NF) v.s. P.A.
@@
Previously, Wu et al. (2020) proposed that one could obtain a more robust model via adversarial weight perturbation. To seek whether such models could also be robust to random bit errors, we conducted an experiment on CIFAR-10 with the proposed adversarially trained PreAct ResNet18. The experimental results are shown in Table 26. We find that the average perturbed accuracy is 23% and 63.2% for PreAct ResNet18 under 1% and 0.5% B.E.R., respectively. This result is lower than 38.9% and 70.1% from ResNet18 in Table 12, indicating their poor generalization ability against random bit errors. Nevertheless, when equipped NeuralFuse on the perturbed model, we could still witness a significant recover percentage under both 1% and 0.5% B.E.R. This result further demonstrates that NeuralFuse could be adapted to various models (i.e., trained in different learning algorithms).
| Base Model | B.E.R. | NF | C.A. | P.A. | C.A. (NF) | P.A. (NF) | R.P. |
@@
[Note] B.E.R.: the bit error rate of the base model, NF: NeuralFuse, C.A. (%): clean accuracy, P.A. (%): perturbed accuracy, and R.P.: total recover percentage of P.A. (NF) v.s. P.A.
@@
To further understand how our proposed NeuralFuse works, we visualize the output distribution from the final linear layer of the base models and project the results onto the 2D space using t-SNE (van der Maaten & Hinton, 2008). Figure 15 shows the output distribution from ResNet18 (trained on CIFAR-10) under a 1% bit error rate. We chose two generators that have similar architecture: ConvL and ConvS, for this experiment. We can observe that: (a) The output distribution of the clean model without NeuralFuse can be grouped into 10 classes denoted by different colors. (b) The output distribution of the perturbed model under a 1% bit error rate without NeuralFuse shows mixed representations and therefore degraded accuracy. (c) The output distribution of the clean model with ConvL shows that applying NeuralFuse will not hurt the prediction of the clean model too much (i.e., it retains high accuracy in the regular voltage setting). (d) The output distribution of the perturbed model with ConvL shows high separability (and therefore high perturbed accuracy) as opposed to (b). (e)/(f) shows the output distribution of the clean/perturbed model with ConvS. For both (e) and (f), we can see nosier clustering when compared to (c) and (d), which means the degraded performance of ConvS compared to ConvL. The visualization validates that NeuralFuse can help retain good data representations under random bit errors and that larger generators in NeuralFuse have better performance than smaller ones.
@@
Clean |
ConvL |
ConvS |
DeConvL |
DeConvS |
UNetL |
@@
Clean |
ConvL |
ConvS |
DeConvL |
DeConvS |
UNetL |

### v3

@@ Appendix I Additional Experiments on Adversarial Training
K-bits CA PA CA (NF) PA (NF) RP 100 92.6 38.9 12.4 92.4 38.3 12.1 -0.6 500 92.1 38.7 12.5 -0.2 5,000 92.6 38.9 12.5 0 20,000 60.1 23.0 8.1 -16 100,000 71.1 23.6 6.6 -16 • Note. CA (): clean accuracy; PA (): perturbed accuracy; NF: NeuralFuse; and RP: total recovery percentage of PA (NF) vs. PA
@@
Previously, Wu et al. (2020) proposed that one could obtain a more robust model via adversarial weight perturbation. To seek whether such models could also be robust to random bit errors, we conducted an experiment on CIFAR-10 with the proposed adversarially trained PreAct ResNet18. The experimental results are shown in Table 26. We find that the average perturbed accuracy is 23% and 63.2% for PreAct ResNet18 under 1% and 0.5% BER, respectively. This result is lower than 38.9% and 70.1% from ResNet18 in Table 12, indicating their poor generalization ability against random bit errors. Nevertheless, when equipped NeuralFuse on the perturbed model, we could still witness a significant recover percentage under both 1% and 0.5% BER This result further demonstrates that NeuralFuse could be adapted to various models (i.e., trained in different learning algorithms).
| Base Model | BER | NF | CA | PA | CA (NF) | PA (NF) | RP |
@@
Note. BER: the bit-error rate of the base model; CA (): clean accuracy; PA (): perturbed accuracy; NF: NeuralFuse; and RP: total recovery percentage of PA (NF) vs. PA
@@
To further understand how our proposed NeuralFuse works, we visualize the output distribution from the final linear layer of the base models and project the results onto the 2D space using t-SNE (van der Maaten and Hinton, 2008). Figure 15 shows the output distribution from ResNet18 (trained on CIFAR-10) under a 1% bit error rate. We chose two generators that have similar architecture: ConvL and ConvS, for this experiment. We can observe that: (a) The output distribution of the clean model without NeuralFuse can be grouped into 10 classes denoted by different colors. (b) The output distribution of the perturbed model under a 1% bit error rate without NeuralFuse shows mixed representations and therefore degraded accuracy. (c) The output distribution of the clean model with ConvL shows that applying NeuralFuse will not hurt the prediction of the clean model too much (i.e., it retains high accuracy in the regular voltage setting). (d) The output distribution of the perturbed model with ConvL shows high separability (and therefore high perturbed accuracy) as opposed to (b). (e)/(f) shows the output distribution of the clean/perturbed model with ConvS. For both (e) and (f), we can see nosier clustering when compared to (c) and (d), which means the degraded performance of ConvS compared to ConvL. The visualization validates that NeuralFuse can help retain good data representations under random bit errors and that larger generators in NeuralFuse have better performance than smaller ones.
@@
| Clean | ||
| ConvL | ||
| ConvS | ||
| DeConvL | ||
| DeConvS | ||
| UNetL | ||
@@
| Clean | |
| ConvL | |
| ConvS | |
| DeConvL | |
| DeConvS | |
| UNetL | |